Shown below is the financial date of company A and company B at end of current year

company A company B

Netsales (All on credit) $ 1440000 $ 1190000

Cost of good sold 1260000 825000

Cash 36000 70000

Accounts receivable (Net) 180000 140000

Inventory 504000 165000

current liabilities 240000 150000

Assume that the year end balance shown for accounts receivables and for inventory also represent the average of those accounts throughout the years.

Required.

from the view point of short term creditors how you assess the the quality of each company working capital to which company you will prefer to sell merchandise of $ 50000 on 30 days open account?


1
Expert's answer
2021-07-14T11:59:50-0400

Solution:

The quality of each company's working capital is measured by comparing the total current assets with the total current liabilities. A company that has more working capital is considered to be in a better financial position.


The working capital calculation for company A:

Current assets = Cash + Accounts receivable + Inventory

Current assets = 36,000 + 180,000 + 504,000 = 720,000

Current liabilities = 240,000

Working capital = Current assets – current liabilities

Working capital = 720,000 – 240,000 = 480,000

 

The working capital calculation for company B:

Current assets = Cash + Accounts receivable + Inventory

Current assets = 70,000 + 140,000 + 165,000 = 375,000

Current liabilities = 150,000

Working capital = Current assets – current liabilities

Working capital = 375,000 – 150,000 = 225,000

 

From the above calculations, company A has the best working capital compared to company B, which means that it has more current assets to cover its current liabilities and still remains with surplus funds for other investment purposes and growing the business.

Therefore, I would prefer to sell merchandise of $50,000 on 30 days open account to company A. This is because company A has a strong working capital, which indicates that it can easily pay off its current liabilities without much effort and time. By selling the merchandise to company A, you are guaranteed that you will be paid on time within 30 days.